Maureen jogged 3 km less than her older sister. If each of them jogged 4 km less, the total distances would have been 15 km. How far did each sister jog?Please, and thank you :)
Answer by ankor@dixie-net.com(15624)

Maureen jogged 3 km less than her older sister.
If each of them jogged 4 km less, the total distances would have been 15 km.
How far did each sister jog
:
let m = distance jogged by M
then
(m+3) = distance jogged by sister
:
"If each of them jogged 4 km less, the total distances would have been 15 km."
m - 4 + m+3 - 4 = 15
2m - 5 = 15
2m = 15 + 5
2m = 20
m = 20/2
m = 10 km jogged by M
then
10+3 = 13 km jogged by sister
:
:
See if that checks out in the statement
"If each of them jogged 4 km less, the total distances would have been 15 km."
10 - 4 = 6 km
13 - 4 = 9 km
--------------
total: 15 km